OpenStudy (anonymous):

Write the product as a trinomial. (m – 5)(m + 2) A. m2 – 3m – 10 B. m2– 3m + 10 C. m2 + 3m – 10 D. m2 – 10

OpenStudy (anonymous):

im thinking C @jojo4eva ???

OpenStudy (anonymous):

Okay, tri = three. In this case D is no applicable since it only has two terms. Now lets use FOIL to obtain the expanded form of \((m – 5)(m + 2)\). FOIL: First Outer Inner Last \[(m – 5)(m + 2) \implies m^2~\text{(First m}\times\text{m)}\] \[(m – 5)(m + 2) \implies 2m~\text{(Outer m}\times\text{2)}\] \[(m – 5)(m + 2) \implies -5m~\text{(Inner -5m}\times\text{m)}\] \[(m – 5)(m + 2) \implies -10~\text{(Last -5}\times\text{2)}\] \[m^2+2m+-5m+(-10) \implies m^2-3 m-10\] \[m^2-3 m-10\]